# A primer on quasi-convex functions in nonlinear potential theories

### v2

@@ 4.3. The Crandall–Ishii–Lions Theorem on Sums
    $\mathcal{F}_{\Delta}:=\{(r,p,A)\in\mathcal{J}^{2}:\ {\rm tr}(A)\geqslant 0\}$
    and its subarmonics are characterterized by
@@
    This subsequation is self-dual $\widetilde{\mathcal{F}}=\mathcal{F}$ in terms of the duality operation recalled in the following section.
@@
    The convexity subequation is $\mathcal{F}:=X\times\mathcal{M}(\mathcal{P})$ where
    $\mathcal{M}(\mathcal{P}):={\mathbb{R}}\times{\mathbb{R}}^{n}\times\mathcal{P}:=\{(r,p,A)\in\mathcal{J}^{2}:\ A\geqslant 0\}$
    and its subarmonics are characterterized by
    $u\in\mathcal{F}(X)\ \Leftrightarrow\ u$ is convex (5.24)
    (away from the connected components on which $u\equiv-\infty$). The subharmonics of the dual subequation are the so-called subaffine functions (they satisfy the comparison principle with respect to all affine functions). This example is the most basic example of a monotonicity cone subequation (a subequation which is also a closed convex cone with vertex at the origin) as will be discussed in the following section. This particular cone is used in the proof of the comparison principle for all pure second order subequations and degenerate elliptic PDEs, as will be discussed in Section 9. A fundamental family of such cones can be found in Chapter 5 of [8].
@@
    Example 5.18 (Quasi-convex functions).
    For $\lambda>0$, The $\lambda$-quasi-convexity subequation is $\mathcal{F}_{\lambda}:=X\times\mathcal{M}(\mathcal{P}_{\lambda})$ where
    $\mathcal{M}(\mathcal{P}_{\lambda}):={\mathbb{R}}\times{\mathbb{R}}^{n}\times\mathcal{P}_{\lambda}:=\{(r,p,A)\in\mathcal{J}^{2}:\ A+\lambda I\geqslant 0\}$
    and its subarmonics are characterterized by
    $u\in\mathcal{F}_{\lambda}(X)\ \Leftrightarrow\ u$ is $\lambda$-quasi-convex (5.25)
    (away from the connected components on which $u\equiv-\infty$). Moreover, it is clear that the notion of being $\mathcal{F}$-subharmonic is local since the notion of upper contact jet is local (see Proposition 7.4(i)). Hence one can characterize the space $\mathsf{qc}_{\text{loc}}(X)$ of locally quasi-convex functions on $X$ as those $u\in\operatorname{USC}(X)$ such that
@@
    6. Duality and monotonicity
    This section is dedicated to two fundamental notions in general second order potential theories. Duality is reviewed first with its important consequence of reformulating $\mathcal{F}$-superharmonics as subharmonics for a dual subequation $\widetilde{\mathcal{F}}$. Next, the unifying notion of monotonicity is reviewed. Together, they give rise to the duality-monotonicity method for proving the validity of the comparison principle, which is a key ingredient in the treatment of uniqueness and uniqueness of solutions to the Dirichlet problem for $\mathcal{F}$-harmonic functions.
    A subequation $\mathcal{F}$ naturally gives rise to a dual subequation $\widetilde{\mathcal{F}}$, which will play a crucial role in all that follows. Consider the following notion of duality, introduced by Harvey and Lawson [23, 24] and is defined in a purely set-theoretic way.
@@ 6. Duality and monotonicity
    It turns out that if $\mathcal{F}$ is a subequation, then $\widetilde{\mathcal{F}}$ is a subequation as well; this is a consequence of the following proposition, which collects some elementary properties of the Dirichlet dual. These properties are to be found in [23, Section 4] in a pure second-order context and in [24, Section 3] for subequations on Riemannian manifolds, as well as in [8, Proposition 3.2] for constant coefficient subequations; we essentially reproduce the proof of this last proposition, in order to point out that it works fine also if one has nonconstant coefficients. This happens because of a consequence of the topological condition (T), as noted in [24]. More precisely, if we define the dual of the fiber $\mathcal{F}_{x}\subset{\mathbb{R}}\times{\mathbb{R}}^{n}\times\mathcal{S}(n)$ relative to the ambient space ×${\mathbb{R}}\times{\mathbb{R}}^{n}\times\mathcal{S}(n)$, one can compute the dual of $\mathcal{F}$ fiberwise; that is,
@@
      $\mathcal{F}=\overline{\operatorname{int}{\mathcal{F}}}$ $\iff$ $\widetilde{\widetilde{\mathcal{F}}}=\mathcal{F}$;
@@
    Properties (4) and (5) follow from (2), and property (8) follows from properties (4), (5) and (7). Property (1) follows from Definition 6.1. For property (2), note that $\mathcal{F}_{x}+J\subset\mathcal{F}_{x}$ implies that $\operatorname{int}{\mathcal{F}}_{x}+J$ is an open subset of $\mathcal{F}$, hence $\operatorname{int}{\mathcal{F}}_{x}+J\subset\operatorname{int}{\mathcal{F}}_{x}$, which yields, taking the complements, ~$\widetilde{\mathcal{F}}_{x}\subset\widetilde{\mathcal{F}}_{x}-J$, as desired. Similarly, since $\operatorname{int}{\mathcal{F}}_{x}+J$ is the interior of $\mathcal{F}_{x}+J$, we have $\widetilde{\mathcal{F}_{x}+J}=\widetilde{\mathcal{F}}_{x}-J$, which is property (3). Note now that by Definition 6.1
    ~~$\widetilde{\widetilde{\mathcal{F}}}=\big{(}\operatorname{int}\big{(}(\operatorname{int}\mathcal{F})^{\mathsf{c}}\big{)}\big{)}^{\mathsf{c}};$ (6.1)
    since for any subset $\mathcal{S}\subset\mathcal{J}^{2}(X)$ one has $\operatorname{int}{\mathcal{S}}={\overline{\mathcal{S}^{\mathsf{c}}}}^{\mathsf{c}}$, equality (6.1) yields ~~$\widetilde{\widetilde{\mathcal{F}}}=\overline{\operatorname{int}{\mathcal{F}}}$. This proves property (6). Similarly, if $\mathcal{F}$ satisfies (T), consider the following chain of equalities:
    $\overline{\operatorname{int}{\widetilde{\mathcal{F}}}}=\overline{{\overline{\widetilde{\mathcal{F}}\,{}^{\mathsf{c}}}\,}^{\mathsf{c}}}=\overline{{\overline{-\operatorname{int}{\mathcal{F}}}\,}^{\mathsf{c}}}=\overline{-\mathcal{F}^{\mathsf{c}}}=(-\operatorname{int}{\mathcal{F}})^{\mathsf{c}}=\widetilde{\mathcal{F}}.$
@@
    By making use of duality, we can reformulate the notions of being $\mathcal{F}$-subharmonic and $\mathcal{F}$-harmonic, which we record in the following observation.
    Remark 6.4 ($\mathcal{F}$-superharmonics by duality).
    For $v\in\operatorname{LSC}(X)$, one can define $v$ to be $\mathcal{F}$-superharmonic on $X$ if $-v$ is $\widetilde{\mathcal{F}}$-subharmonic on $X$. This is equivalent to 5.9; indeed, the condition $(v(x),q,B)\notin\operatorname{int}{\mathcal{F}}_{x}$ for each $(q,B)\in J^{2,-}_{x}v$ is equivalent to $(-v(x),p,A)\notin-\operatorname{int}{\mathcal{F}}_{x}$ for each $(p,A)=(-q,-B)\in J^{2,+}_{x}(-v)$, since it is immediate to see (cf. Definition 2.6) that $-J^{2,-}_{x}v=J^{2,+}_{x}(-v)$. This means that $v$ is $\mathcal{F}$-superharmonic at $x$ if and only if $-v$ is $\widetilde{\mathcal{F}}$-subharmonic at $x$, because
@@
    Finally, this yields the following equivalent definition of viscosity solution: a function $u\in C(X)$ is $\mathcal{F}$-harmonic on $X$ if and only if $u\in\mathcal{F}(X)$ and $-u\in\widetilde{\mathcal{F}}(X)$.
@@
    $\mathcal{F}_{x}+\mathcal{M}_{x}\subset\mathcal{F}_{x}\ \ \text{for each}\ x\in X.$ (6.2)
@@
    First and foremost, the properties (P) and (N) are monotonicity properties. As noted previously, property (P) for subequations $\mathcal{F}$ corresponds to degenerate elliptic operators $F$ and properties (P) and (N) together correspond to proper elliptic operators. Properties (P) plus (N) can be expressed as the single monotonicity property
    $\mathcal{F}_{x}+\mathcal{M}_{0}\subset\mathcal{F}_{x}\ \ \text{for each}\ x\in X$ (6.3)
@@
    $\mathcal{M}_{0}:=\mathcal{N}\times\{0\}\times\mathcal{P}\subset\mathcal{J}^{2}={\mathbb{R}}\times{\mathbb{R}}^{n}\times\mathcal{S}(n)$ (6.4)
@@
    $\mathcal{N}:=\{r\in{\mathbb{R}}:\ r\leqslant 0\}\quad\text{and}\quad\mathcal{P}:=\{P\in\mathcal{S}(n):\ P\geqslant 0\}.$ (6.5)
    Hence $\mathcal{M}_{0}$ will be referred to as the minimal monotonicity cone in $\mathcal{J}^{2}$. However, it is important to remember that $\mathcal{M}_{0}\subset\mathcal{J}^{2}$ is not a subequation since it has empty interior so that property (T) fails.
    Monotonicity is also used to formulate reductions when certain jet variables are “silent” in the subequation constraint $\mathcal{F}$. See section 2.2 of [32] for details.
    Most importantly, monotonicity when combined with duality and fiberegularity in the variable coefficient case (the natural continuity of the fiber map $x\mapsto\mathcal{F}_{x}$)), yields has a very general, flexible and elegant geometrical approach to comparison when a subequation $\mathcal{F}$ admits a constant coefficient monotonicity cone subequation $\mathcal{M}$. This will be described section 9.
    7. Basic tools in nonlinear potential theory
    In this section, we introduce various results that are included in the “basic tool kit of viscosity solution techniques” in [8]: the Bad Test Jet Lemma 7.1, the Definitional Comparison Lemma 7.2 and Proposition 7.4 which gathers various widely used constructions of $\mathcal{F}$-harmarmonics obatined by alegebraic opertions, optimization procedures and limits. We begin with the first tool which is very useful when one seeks to check the validity of subharmonicity at a point by a contradiction argument. More precisely, if $u$ fails to be subharmonic at a given point, then one must have the existence of a bad test jet at that point, as stated in the following lemma. This criterion is essentially the contrapositive of the definition of viscosity subsolution, when one takes strict upper contact quadratic functions as upper test functions (see [8, Lemma 2.8 and Lemma C.1]). Nevertheless, for the benefit of the reader, we provide a brief proof based on LABEL:defn:FSH ($\mathcal{F}$-subharmonics).
@@ 7. Basic tools in nonlinear potential theory
    Given $u\in\operatorname{USC}(X)$, $x\in X$ and $\mathcal{F}_{x}\neq\emptyset$, suppose $u$ is not $\mathcal{F}$-subharmonic at $x$. Then there exists $\varepsilon>0$ and a $2$-jet $J\notin\mathcal{F}_{x}$ such that the (unique) quadratic function $\varphi_{J}$ with $J^{2}_{x}\varphi_{J}=J$ is an upper test function for $u$ at $x$ in the following $\varepsilon$-strict sense:
@@
    If $u$ fails to be $\mathcal{F}$-subharmonic at $x$, then according to LABEL:defn:FSH there exists an upper test function $\varphi$ for $u$ at $x$ such that
    $J:=J^{2}\varphi_{x}\not\in\mathcal{F}_{x}$
    Since the complement of $\mathcal{F}_{x}$ (with respect to ×${\mathbb{R}}\times{\mathbb{R}}^{n}\times\mathcal{S}(n)$) is open, we have that $J^{\prime}\vcentcolon=J+(0,0,2\varepsilon I)\notin\mathcal{F}_{x}$ for $\varepsilon$ small. Also, $J^{\prime}$ is a quadratic upper test jet for $u$ in $x$ by Definition 2.6; that is, if $(p,A)$ is an upper contact jet for $u$ at $x$, so is $(p,A+\eta I)$ for any $\eta>0$. This means that
    $u(y)\leqslant\varphi_{J^{\prime}}(y)\quad\text{$\forall y$ near $x$ (with equality at $x$)},$
    but $\varphi_{J^{\prime}}=\varphi_{J}-\varepsilon|\cdot-\,x|^{2}$, hence (7.1) follows. ∎
    The second tool is a comparison principle whose validity characterizes the $\mathcal{F}$-subharmonic functions for a given subequation $\mathcal{F}$. We begin by recalling two equivalent forms of potential theoretic comparison which will play an important role in all of our proofs of comparison principles.
@@
    for all $u\in\operatorname{USC}(\overline{\Omega})$, $w\in\operatorname{LSC}(\overline{\Omega})$ which are $\mathcal{F}$-subharmonic, $\mathcal{F}$-superharmonic on $\Omega$, respectively. Making use of the duality reformulation with $v\vcentcolon=-w$, comparison (CP) is equivalent to
    $u+v\leqslant 0\ \text{on $\partial\Omega$}\implies u+v\leqslant 0\ \text{on $\Omega$}$ (${}^{\prime}$)
    for all $u,v\in\operatorname{USC}(\overline{\Omega})$ which are, respectively, $\mathcal{F}$-, $\widetilde{\mathcal{F}}$-subharmonic on $\Omega$. This in turn is equivalent to the following zero maximum principle:
@@
    for all $z\in\operatorname{USC}(\overline{\Omega})\cap\big{(}\mathcal{F}(\Omega)+\widetilde{\mathcal{F}}(\Omega)\big{)}$.
    We are now ready for the lemma, which states that comparison holds if the function $z$ in (ZMP) is the sum of a $\mathcal{F}$-subharmonic and a $C^{2}$-smooth and strictly $\widetilde{\mathcal{F}}$-subharmonic. It is called definitional comparison because it relies only upon the “good” definitions we gave for subharmonic and for subequations, which includes the negativity condition (N) which is important in the proof. It was stated and proven by Cirant, Harvey, Lawson, and Payne [8, Lemma 3.14] in a context of constant coefficient subequations; we reproduce here its proof to highlight that it holds in the nonconstant coefficient case, too.
@@
      If $u$ is $\mathcal{F}$-subharmonic on $X$, then the following form of the comparison principle holds for each bounded domain $\Omega\Subset X$:
@@
      Conversely, suppose that for each $x\in X$ there is a neighborhood $\Omega\Subset X$ of $x$ where the above form of comparison holds. Then $u$ is $\mathcal{F}$-subharmonic on $X$.
@@
    If the form of comparison in (i) fails for some domain $\Omega\Subset X$ and some regular strict $\widetilde{\mathcal{F}}$-subharmonic function $v$, then $u+v\in\operatorname{USC}(\overline{\Omega})$ will have a positive maximum value $m>0$ at an interior point $x_{0}\in\Omega$ and hence $\varphi\vcentcolon=-v+m$ is $C^{2}$ near $x_{0}$ and satisfies
    $u-\varphi\leqslant 0\quad\text{near $x_{0}$, with equality at $x_{0}$}.$
    Since $u$ is $\mathcal{F}$-subharmonic at $x_{0}$, this implies that
@@
    which contradicts the property (N) for $\mathcal{F}$ since $m>0$ and
    $J^{2}_{x_{0}}v\in\operatorname{int}{\widetilde{\mathcal{F}}}_{\!x_{0}}=(-\mathcal{F}_{x_{0}})^{\mathsf{c}}.$
    Conversely, suppose that $u$ fails to be $\mathcal{F}$-subharmonic at some $x_{0}\in X$. By the Bad Test Jet Lemma 7.1, there exist $\rho,\varepsilon>0$ and an upper contact jet $(p,A)\in J^{2,+}_{x_{0}}u$ such that $J=(u(x_{0}),p,A)\notin\mathcal{F}_{x_{0}}$ and
@@
    Sometimes it is useful to prove the contrapositive of the form of comparison in part (i) of Lemma 7.2 in order to conclude $\mathcal{F}$-subharmonicity. That is to say, in order to show by (ii) that $u$ is $\mathcal{F}$-subharmonic on $X$ one proves that, for each $x\in X$ there is a neighborhood $\Omega\Subset X$ of $x$ where
@@
    for every $v\in\operatorname{USC}(\overline{\Omega})\cap C^{2}(\Omega)$ which is strictly $\widetilde{\mathcal{F}}$-subharmonic on $\Omega$. Conversely, one can also infer that the implication (7.3) holds whenever one knows that $u$ is $\mathcal{F}$-subharmonic on $X$. In situations where we are interested in proving the $\mathcal{F}$-subharmonicity of a function which is somehow related to a given $\mathcal{F}$-subharmonic, this helps to close the circle (for example, see the proof of upcoming Proposition 7.4).
    Finally, we give a collection of tools which represents elementary properties shared by functions in $\mathcal{F}(X)$, the set of $\mathcal{F}$-subharmonics on $X$. They are to be found in [23, Section 4] for pure second-order subequations, in [24, Theorem 2.6] for subequations on Riemannian manifolds, in [8, Proposition D.1] for constant-coefficient subequations. We place it here because by invoking the definitional comparison lemma one can perform most of the proofs along the lines of those in Harvey–Lawson [23]. Indeed, we are going to use the definitional comparison in order to make up for the lack, for arbitrary subequations, of a result like [23, Lemma 4.6].

### v3

@@ 5.2. Some simple classical examples
$\mathcal{F}_{\Delta}\vcentcolon=\{(r,p,A)\in\mathcal{J}^{2}:\ {\rm tr}(A)\geqslant 0\}$
and its subarmonics are characterized by
@@
This subequation is self-dual ($\widetilde{\mathcal{F}}=\mathcal{F}$) in terms of the duality operation discussed in the following chapter.
@@
The convexity subequation is $\mathcal{F}=X\times\mathcal{M}(\mathcal{P})$ where
$\mathcal{M}(\mathcal{P})\vcentcolon={\mathbb{R}}\times{\mathbb{R}}^{n}\times\mathcal{P}\vcentcolon=\{(r,p,A)\in\mathcal{J}^{2}:\ A\geqslant 0\}$
and its subharmonics are characterized by
$u\in\mathcal{F}(X)\ \Leftrightarrow\ u$ is locally convex
(away from the connected components on which $u\equiv-\infty$). In particular, if $X$ is also convex and $u$ is real valued then $u\in\mathcal{F}(X)$ if and only if $u$ is convex on $X$. On the other hand, the subharmonics of the dual subequation are the so-called subaffine functions; that is, they satisfy the comparison principle with respect to all affine functions. The characterizations of these two fundamental examples of constant coefficient pure second order suabequations will be briefly discussed in Section 9.2 (see Proposition 9.5).
The convexity subequation is the most basic example of a monotonicity cone subequation (a constant coefficient subequation which is also a closed convex cone with vertex at the origin; see Section 6.2). This particular cone is used in the proof of the comparison principle for all pure second order subequations and degenerate elliptic PDEs, as will be discussed in Chapter 9. A fundamental family of such cones can be found in [9, Chapter 5]; see also Remark 9.16.
@@
Example 5.18 (Semiconvex functions).
For $\lambda>0$, the $\lambda$-semiconvexity subequation is $\mathcal{F}_{\lambda}=X\times\mathcal{M}(\mathcal{P}_{\lambda})$ where
$\mathcal{M}(\mathcal{P}_{\lambda})\vcentcolon={\mathbb{R}}\times{\mathbb{R}}^{n}\times\mathcal{P}_{\lambda}\vcentcolon=\{(r,p,A)\in\mathcal{J}^{2}:\ A+\lambda I\geqslant 0\}$
and its subharmonics are characterized by
$u\in\mathcal{F}_{\lambda}(X)\ \Leftrightarrow\ u$ is locally $\lambda$-semiconvex
(away from the connected components on which $u\equiv-\infty$). In particular, if $X$ is also convex and $u$ is real valued then $u\in\mathcal{F}_{\lambda}(X)$ if and only if $u$ is $\lambda$-semiconvex on $X$.
[…]
@@
Chapter 6 Duality and monotonicity in general potential theories
[…]
6.1. Duality
A subequation $\mathcal{F}$ naturally gives rise to a dual subequation $\widetilde{\mathcal{F}}$, which will play a crucial role in all that follows. Consider the following notion of duality, introduced by Harvey and Lawson [26, 27] and defined in a purely set-theoretic way.
@@ 6.1. Duality
It turns out that if $\mathcal{F}$ is a subequation, then $\widetilde{\mathcal{F}}$ is a subequation as well; this is a consequence of the following proposition, which collects some elementary properties of the Dirichlet dual. These properties are to be found in [26, Section 4] in a pure second-order context and in [27, Section 3] for subequations on Riemannian manifolds, as well as in [9, Proposition 3.2] for constant coefficient subequations. We essentially reproduce the proof of this last proposition in order to stress that it also works well if one has non-constant coefficients. This happens because of a consequence of the topological condition (T), as noted in [27]: if we define the dual of the fiber $\mathcal{F}_{x}\subset{\mathbb{R}}\times{\mathbb{R}}^{n}\times\mathcal{S}(n)$ relative to the ambient space ×${\mathbb{R}}\times{\mathbb{R}}^{n}\times\mathcal{S}(n)$, one can compute the dual of $\mathcal{F}$ fiberwise; that is,
@@
    $\mathcal{F}=\overline{\operatorname{int}{\mathcal{F}}}$ $\iff$ $\hskip 2.5pt\widetilde{\widetilde{\mathcal{F}}}=\mathcal{F}$;
@@
Property (1) follows from Definition 6.1. Properties (4) and (5) follow from (2), and property (8) follows from properties (4), (5) and (7).
For property (2), note that $\mathcal{F}_{x}+J\subset\mathcal{F}_{x}$ implies that $\operatorname{int}{\mathcal{F}}_{x}+J$ is an open subset of $\mathcal{F}$, hence $\operatorname{int}{\mathcal{F}}_{x}+J\subset\operatorname{int}{\mathcal{F}}_{x}$, which yields, taking the complements, ~$\widetilde{\mathcal{F}}_{x}\subset\widetilde{\mathcal{F}}_{x}-J$, as desired. Similarly, since $\operatorname{int}{\mathcal{F}}_{x}+J$ is the interior of $\mathcal{F}_{x}+J$, we have $\widetilde{\mathcal{F}_{x}+J}=\widetilde{\mathcal{F}}_{x}-J$, which is property (3).
Note now that by Definition 6.1
~~$\hskip 2.5pt\widetilde{\widetilde{\mathcal{F}}}=\Bigl{(}\operatorname{int}\bigl{(}(\operatorname{int}\mathcal{F})^{\mathsf{c}}\bigr{)}\Bigr{)}^{\mathsf{c}};$ (6.1)
since for any subset $\mathcal{S}\subset\mathcal{J}^{2}(X)$ one has $\operatorname{int}{\mathcal{S}}={\overline{\mathcal{S}^{\mathsf{c}}}\,}^{\mathsf{c}}$, equality (6.1) yields ~~$\hskip 2.5pt\widetilde{\widetilde{\mathcal{F}}}=\overline{\operatorname{int}{\mathcal{F}}}$, thus proving property (6). Similarly, if $\mathcal{F}$ satisfies (T), consider the following chain of equalities:
$\overline{\operatorname{int}{\widetilde{\mathcal{F}}}}=\overline{{\overline{\widetilde{\mathcal{F}}\,{\vphantom{I}}^{\mathsf{c}}}\,}^{\mathsf{c}}}=\overline{{\overline{-\operatorname{int}{\mathcal{F}}}\,}^{\mathsf{c}}}=\overline{-\mathcal{F}^{\mathsf{c}}}=(-\operatorname{int}{\mathcal{F}})^{\mathsf{c}}=\widetilde{\mathcal{F}};$
@@
[…]
$(v(x),q,B)\notin\operatorname{int}{\mathcal{F}}_{x}$  for each $(q,B)\in J^{2,-}_{x}v$,
which is equivalent to
$(-v(x),p,A)\notin-\operatorname{int}{\mathcal{F}}_{x}$  for each $(p,A)=(-q,-B)\in J^{2,+}_{x}(-v)$,
[…]
@@
[…]
6.2. Monotonicity
@@
$\mathcal{F}_{x}+\mathcal{M}_{x}\subset\mathcal{F}_{x}\quad\forall x\in X.$ (6.2)
@@
First and foremost, the properties (P) and (N) are monotonicity properties. As noted previously, property (P) for subequations $\mathcal{F}$ corresponds to degenerate elliptic operators $F$ and properties (P) and (N) together correspond to proper elliptic operators. They can be expressed as the single monotonicity property
$\mathcal{F}_{x}+\mathcal{M}_{0}\subset\mathcal{F}_{x}\quad\forall x\in X$
@@
$\mathcal{M}_{0}\vcentcolon=\mathcal{N}\times\{0\}\times\mathcal{P}\subset\mathcal{J}^{2}={\mathbb{R}}\times{\mathbb{R}}^{n}\times\mathcal{S}(n)$ (6.3)
@@
$\mathcal{N}\vcentcolon=\{r\in{\mathbb{R}}:\ r\leqslant 0\}\quad\text{and}\quad\mathcal{P}\vcentcolon=\{P\in\mathcal{S}(n):\ P\geqslant 0\}.$ (6.4)
Hence $\mathcal{M}_{0}$ will be referred to as the minimal monotonicity cone in $\mathcal{J}^{2}$. However, it is important to remember that $\mathcal{M}_{0}\subset\mathcal{J}^{2}$ is not a subequation since it has empty interior, so that property (T) fails.
Monotonicity is also used to formulate reductions when certain jet variables are “silent” in the subequation constraint $\mathcal{F}$. See [35, Section 2.2] for details, as well as the examples in Section 9.2.
Most importantly, monotonicity, when combined with duality and fiberegularity in the variable coefficient case (the natural continuity of the fiber map $x\mapsto\mathcal{F}_{x}$), yields a very general, flexible and elegant geometrical approach to comparison when a subequation $\mathcal{F}$ admits a constant coefficient monotonicity cone subequation $\mathcal{M}$. This will be described in Chapter 9.
Chapter 7 Basic tools in nonlinear potential theory
[…]
7.1. The Bad Test Jet and Definitional Comparison Lemmas
We begin with the first useful tool which allows one to check the validity of subharmonicity at a point by a contradiction argument. More precisely, if $u$ fails to be subharmonic at a given point, then one must have the existence of a bad test jet at that point, as stated in the following lemma. This criterion is essentially the contrapositive of the definition of viscosity subsolution, when one takes strict upper contact quadratic functions as upper test functions (see [9, Lemmas 2.8 and C.1]). Nevertheless, for the benefit of the reader, we provide a brief proof based on Definition 5.1.
@@ 7.1. The Bad Test Jet and Definitional Comparison Lemmas
[…]
@@
[…]
Definition 7.2.
@@
[…]
Making use of the duality reformulation with $v\vcentcolon=-w$, comparison (CP) is equivalent to
$u+v\leqslant 0\ \text{on $\partial\Omega$}\implies u+v\leqslant 0\ \text{on $\Omega$}$ (CP)
[…]
@@
for all $z\in\operatorname{USC}(\overline{\Omega})\cap\bigl{(}\mathcal{F}(\Omega)+\widetilde{\mathcal{F}}(\Omega)\bigr{)}$.
[…]
@@
[…]
@@
[…]
@@
which contradicts the property (N) for $\mathcal{F}$ since $m>0$ and $J^{2}_{x_{0}}v\in\operatorname{int}{\widetilde{\mathcal{F}}}_{\!x_{0}}=(-\mathcal{F}_{x_{0}})^{\mathsf{c}}$.
[…]
@@
[…]
@@
[…]
7.2. Fundamental properties of subharmonics
[…]